OpenStudy (anonymous):

A sealed room in a hospital, measuring 2 m wide, 7 m long, and 4 m high, is filled with pure oxygen. One cubic meter contains 1000 L, and 22.4 L of any gas contains 6.02 * 10^23 molecules (Avogadro's number). How many molecules of oxygen are there in the room? (Round your answer to two decimal places.)

OpenStudy (anonymous):

I hate Chemistry really...but this is Chemistry Math so I'll try it :P Info : 2 x 7 x 4 dimesion 1 m^3= 1000L 22.4L -----6.02 x10^23 molecules So have to find volume first: V= W x L x H = 2 x 7 x 4 = 56 m^3 56m^3 = 56 x 1000 L = 56000 L so now : 22.4 L ---6.02 x10^23 moles 56000L ---? moles Molecules = 56000L x 6.02x10^23/ 22.4= 1.51 x 10^27 moles